### v1

@@ 3 Discretizing Constraints
The natural discretization
[…]
correctly approximates $\mathcal{C}$ when we scale the entries of $x_{1}$ by the grid size:
$x_{1}[i]=f(t_{1}[i])(\Delta t)^{1/p}.$
With this scaling, the discrete norm approximates the continuous norm:
$c^{p}\coloneqq\left\lVert x_{1}\right\rVert_{p}^{p}=\sum_{i=1}^{I_{1}}f(t_{1}[i])^{p}\Delta t\approx\int_{\mathcal{D}}f(t)^{p}\,dt=\left\lVert f\right\rVert_{p}^{p},$
as $I_{1}\to\infty$.
At coarse scale $s$, we modify the constraint to
@@
$\frac{I_{1}}{I_{s}}\sum_{i=1}^{I_{s}}(x_{s}[i])^{p}\approx\sum_{i=1}^{I_{1}}(x_{1}[i])^{p},$
which gives
$c^{p}=\frac{I_{1}}{I_{s}}\left\lVert x_{s}\right\rVert_{p}^{p}\approx\left\lVert x_{1}\right\rVert_{p}^{p}.$
@@
where
$A_{s}[k,i]=g_{k}(t_{s}[i])(\Delta t)^{1/2}\quad\text{and}\quad x_{s}[i]=f(t_{s}[i])(\Delta t)^{1/2}.$
We bound the discretization error to justify our constraint modification across scales. We assume grid points double at each scale refinement ($I_{s}/I_{s+1}=2$) and use a uniform grid $\{t_{i}\}$ on the domain $\mathcal{D}=[\ell,u]$ to sample an $L_{f}$-Lipschitz function $f$ according to $x[i]=f(t_{i})$. The following proposition quantifies how linear constraints transform under coarsening.
@@
Let $a\in\mathbb{R}^{I}$ discretize a bounded $L_{g}$-Lipschitz function $g:[\ell,u]\to\mathbb{R}$, and let $x$ discretize a bounded $L_{f}$-Lipschitz function $f$. Define the Lipschitz constant for the product function $fg$ as $L_{fg}=L_{f}\left\lVert g\right\rVert_{\infty}+L_{g}\left\lVert f\right\rVert_{\infty}$.
If $\langle a,x\rangle=b$, then the coarsened constraint satisfies the bounds
$\left\lvert\left\langle\overline{a},\overline{x}\right\rangle-\tfrac{1}{2}b\right\rvert$ $\leq\tfrac{1}{4}L_{fg}(u-l)\left(\frac{I}{I-1}\right)$ $\quad\text{for even $I$},$
$\left\lvert\left\langle\overline{a},\overline{x}\right\rangle-\tfrac{1}{2}\frac{I+1}{I}b\right\rvert$ $\leq\tfrac{1}{2}L_{fg}(u-l)$ $\quad\text{for odd $I$}.$
Proof 3.2.
See LABEL:sec-linear-constraint-scaling-proof.
Proposition˜3.1 extends naturally to multiple linear constraints, as described in the following corollary.
@@
Let $g_{k}:[\ell,u]\to\mathbb{R}$ be bounded $L_{g_{k}}$-Lipschitz functions for $k\in[K]$. Let $A\in\mathbb{R}^{K\times I}$ be the discretization $A[k,i]=g_{k}(t[i])$, and define $\overline{A}=A[:,\texttt{begin}\mathbin{:}2\mathbin{:}\texttt{end}]$ as the $(K\times\left\lceil I\right\rceil_{e}/2)$-submatrix with every other column removed, where $\left\lceil I\right\rceil_{e}$ rounds up to the nearest even integer. Define $L_{fg}=\max_{k\in[K]}(L_{fg_{k}})=\max_{k\in[K]}(L_{f}\left\lVert g_{k}\right\rVert_{\infty}+L_{g_{k}}\left\lVert f\right\rVert_{\infty})$ as the maximum Lipschitz constant of the product functions $\{fg_{k}\}_{k\in[K]}$.
If $Ax=b\in\mathbb{R}^{K}$, then
$\left\lVert\overline{A}\overline{x}-\tfrac{1}{2}b\right\rVert_{2}$ $\leq\sqrt{K}\cdot\tfrac{1}{4}L_{fg}(u-l)\left(\frac{I}{I-1}\right)$ $\quad\text{for even $I$},$
$\left\lVert\overline{A}\overline{x}-\frac{I+1}{I}\tfrac{1}{2}b\right\rVert_{2}$ $\leq\sqrt{K}\cdot\tfrac{1}{2}L_{fg}(u-l)$ $\quad\text{for odd $I$}.$
Proof 3.4.
See LABEL:sec-matrix-linear-constraint-proof.
@@
$\left\lvert\left\lVert\overline{x}\right\rVert_{1}-\tfrac{1}{2}b\right\rvert$ $\leq\tfrac{1}{4}L_{f}(u-l)\left(\frac{I}{I-1}\right)$ $\quad\text{for even $I$},$
$\left\lvert\left\lVert\overline{x}\right\rVert_{1}-\frac{I+1}{I}\tfrac{1}{2}b\right\rvert$ $\leq\tfrac{1}{2}L_{f}(u-l)$ $\quad\text{for odd $I$}.$
Proof 3.6.
The proof is similar to the proof of Proposition˜3.1 (see LABEL:sec-linear-constraint-scaling-proof) by considering the vector $a=\mathrm{sign}(x)$, and using the Lipschitz constant $L_{f}$ for $f$ to bound $\left\lvert x[i]-x[i+1]\right\rvert$ in place of $L_{fg}$.
@@
$\left\lvert\left\lVert\overline{z}\right\rVert_{1}-\tfrac{1}{2}\right\rvert\leq\tfrac{1}{4b}L_{f}(u-l)\left(\frac{I}{I-1}\right).$
@@ 4 Convergence of a Multiscale Method
Both the standard and multiscale approaches execute many iterations at the finest scale and converge because, by assumption, they both apply a convergent algorithm at the finest scale. The multiscale achieves faster convergence by warm starting the fine-scale iterations with solutions from coarser scales. The next subsections develop this analysis.
@@ 4.1 Functional Analysis Lemmas
See LABEL:sec-lipschitz-interpolation-proof.
Figure 6: Example for Lemma˜4.1 with the function $x=f(t)=t-\cos(3\pi t)$ on the interval $\ell,u]=[0,1]$. The function $f$ must lie within the dashed parallelogram since it is $L_{f}=1+3\pi$ Lipschitz. Lemma˜4.1 uses the parallelogram constraint to bound the distance between $f$ and the linear interpolation $\lambda_{1}f(a)+\lambda_{2}f(b)$. Lemma˜4.3 shows that the bound in Eq.˜6 is tight for any given $\lambda_{1},\lambda_{2}\geq 0$ with $\lambda_{1}+\lambda_{2}=1$.
@@
Given $\lambda_{1},\lambda_{2}\geq 0$ with $\lambda_{1}+\lambda_{2}=1$, an interval $\ell,u]$, and constant $L_{f}>0$, there exists an $L_{f}$-Lipschitz function on $\ell,u]$ such that Eq.˜6 holds with equality.
@@
Construct
$f(t)=L_{f}\left\lvert t-(\lambda_{1}\ell+\lambda_{2}u)\right\rvert$.
This function is $L_{f}$-Lipschitz with end-point values $f(\ell)=L_{f}\lambda_{2}\left\lvert\ell-u\right\rvert$ and $f(u)=L_{f}\lambda_{1}\left\lvert\ell-u\right\rvert$, and $f(\lambda_{1}\ell+\lambda_{2}u)=0$. Therefore,
$\left\lvert\lambda_{1}f(\ell)+\lambda_{2}f(u)-f(\lambda_{1}\ell+\lambda_{2}u)\right\rvert$
$\quad\qquad=\big\lvert\lambda_{1}L_{f}\lambda_{2}\left\lvert\ell-u\right\rvert+\lambda_{2}L_{f}\lambda_{1}\left\lvert\ell-u\right\rvert\big\rvert$
$\quad\qquad=2L_{f}\lambda_{1}\lambda_{2}\left\lvert\ell-u\right\rvert$.
To bound the error introduced when interpolating a coarse-scale solution to a finer scale, we apply Lemma˜4.1 repeatedly with centre point interpolation ($\lambda_{1}=\lambda_{2}=1/2$). This bounds the error between an exact fine-scale discretization $x_{s}^{*}[j]=f(t_{s}[j])$ and the linear interpolation $x_{s}=\underline{x}_{s+1}$ obtained from a coarser discretization $x_{s+1}[i]=f(t_{s+1}[i])$. See Fig.˜7 (top).
@@
Let $f:[\ell,u]\to\mathbb{R}$ be $L_{f}$-Lipschitz continuous. On the interval $\ell,u]$, let $t_{s}$ be a fine grid with J=2I1J=2I-1 points, and $t_{s+1}$ be a coarse grid with $I$ points. Discretize the function exactly at both scales to obtain vectors $x^{*}_{s}\in\mathbb{R}^{J}$ and $x^{*}_{s+1}\in\mathbb{R}^{I}$ where $x^{*}_{s}[j]=f(t_{s}[j])$ and $x^{*}_{s+1}[i]=f(t_{s+1}[i])$. Let $x_{s+1}=x^{*}_{s+1}\in\mathbb{R}^{I}$ and linearly interpolate to obtain $x_{s}=\underline{x}_{s+1}\in\mathbb{R}^{J}$ where
$\underline{x}_{s+1}[j]=\begin{cases}x_{s+1}\left[\frac{j+1}{2}\right]&\text{if $j$ is odd,}\\ \frac{1}{2}\left(x_{s+1}\left[\frac{j}{2}\right]+x_{s+1}\left[\frac{j}{2}+1\right]\right)&\text{if $j$ is even.}\end{cases}$
Then the difference between the interpolated $x_{s}$ and exact values $x_{s}^{*}$ is bounded by
(7) $\left\lVert x_{s}-x_{s}^{*}\right\rVert_{2}\leq L_{f}\frac{1}{2\sqrt{I-1}}\left\lvert u-\ell\right\rvert$.}
@@
See LABEL:sec-exact-interpolation-proof.
When the coarse discretization contains error, the interpolation analysis extends naturally. Suppose we interpolate not from exact values $x_{s+1}[i]=x_{s+1}^{*}[i]=f(t_{s+1}[i])$, but from approximate values $x_{s+1}[i]=x_{s+1}^{*}[i]+\delta[i]=f(t_{s+1}[i])+\delta[i]$, where $\delta$ is an $I$-vector that represents the error at the coarse scale. Fig.˜7 (bottom) visualizes this scenario.
@@
(8) $\left\lVert x_{s}-x_{s}^{*}\right\rVert_{2}\leq\sqrt{2}\left\lVert\delta\right\rVert_{2}+L_{f}\frac{1}{2\sqrt{I-1}}\left\lvert u-\ell\right\rvert$.}
@@
Figure 7: Illustration of Lemma˜4.5 (top) and Lemma˜4.7 (bottom) for the function $f(t)=72t^{3}-80t^{2}+10t+1$. Both lemmas bound the error between the true fine-scale discretization $x^{*}_{s}$ (orange circles) and the linear interpolation $x_{s}=\underline{x}_{s+1}$ (green squares) obtained from a coarser-scale discretization. The top figure shows Eq.˜7, where $x_{s+1}=x^{*}_{s+1}$ is exact. The bottom figure shows Eq.˜8, where $x_{s+1}=x^{*}_{s+1}+\delta$ contains error.
@@ 4.2 Greedy Multiscale
We use the following approach to show convergence of the greedy multiscale algorithm described in Algorithm˜2. At each scale $s$, we use the iterate convergence (Definition˜2.6) to bound the error between our iterate $x_{s}^{K_{s}}$ and the solution $x_{s}^{*}$ at that scale. Then we use the inexact interpolation lemma (Lemma˜4.7) to bound the error between our iterate $x_{s}^{K_{s}}$ and interpolation $\underline{x}_{S}^{K_{S}}$. This interpolated iterate becomes the initialization $x_{S-1}^{0}=\underline{x}_{S}^{K_{S}}$ at the next finest scale.
@@
Let $q$ denote the rate of convergence of an algorithm applied to a loss function $\tilde{\mathcal{L}}$, let $L_{f^{*}}$ be the Lipschitz constant of the solution function $f^{*}$, and let $S$ denote the coarsest scale. The greedy multiscale method in Algorithm˜2 returns iterate $x_{1}^{K_{1}}$ at the finest scale satisfying
$\left\lVert x^{K_{1}}_{1}-x^{*}_{1}\right\rVert_{2}\leq\sqrt{2^{S-1}}(q)^{r_{S}}\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert_{2}+L_{f^{*}}\frac{\left\lvert u-\ell\right\rvert}{2\sqrt{2^{S+1}}}\sum_{s=1}^{S-1}2^{s}(q)^{r_{s}},$
@@
See LABEL:sec-greedy-multiscale-descent-error-proof.

### v2

@@ 3 Discretizing Constraints
The natural discretization $\tilde{\mathcal{C}}_{1}=\{x_{1}\in\mathbb{R}^{I_{1}}|\left\lVert x_{1}\right\rVert_{p}=c\}$ correctly approximates $\mathcal{C}$ when we scale the entries of $x_{1}$ by the grid size: $x_{1}[i]=f(t_{1}[i])(\Delta t)^{1/p}.$ With this scaling, the discrete norm approximates the continuous norm:
$c^{p}=\left\lVert x_{1}\right\rVert_{p}^{p}=\textstyle\sum_{i=1}^{I_{1}}f(t_{1}[i])^{p}\Delta t\approx\int_{\mathcal{D}}f(t)^{p}\,dt=\left\lVert f\right\rVert_{p}^{p},$
as $I_{1}\to\infty$. At coarse scale $s$, we modify the constraint to
@@
$c^{p}=\frac{I_{1}}{I_{s}}\left\lVert x_{s}\right\rVert_{p}^{p}=\frac{I_{1}}{I_{s}}\sum_{i=1}^{I_{s}}(x_{s}[i])^{p}\approx\sum_{i=1}^{I_{1}}(x_{1}[i])^{p}=\left\lVert x_{1}\right\rVert_{p}^{p}.$
@@
where $A_{s}[k,i]=g_{k}(t_{s}[i])(\Delta t)^{1/2}\ \text{and}\ x_{s}[i]=f(t_{s}[i])(\Delta t)^{1/2}.$
We bound the discretization error to justify our constraint modification across scales. To do this, we extend of the definition of Lipschitz continuity to vectors.
Definition 3.1 (Lipschitz Vector).
A vector $x\in\mathbb{R}^{I}$ is $L_{x}$-Lipschitz when,
$\left\lvert x[i]-x[j]\right\rvert\leq L_{x}\left\lvert i-j\right\rvert\qquad\text{for all}\qquad i,j\in[I].$
We also get a correspondence between Lipschitz functions and vectors.
Corollary 3.2 (Lipschitz Vectors and Functions).
Given $I$ samples $x[i]=f(t[i])$ of an $L_{f}$-Lipschitz function $f:\mathbb{R}\to\mathbb{R}$ on a uniformly spaced grid $\{t[i]\}$, the vector $x\in\mathbb{R}^{I}$ is $L_{x}=L_{f}\Delta t$ Lipschitz.
Proof 3.3.
Let $i,j\in[I]$. Our grid has points $t[i]=l+(i-1)\Delta t$, so we have
$\left\lvert x[i]-x[j]\right\rvert$ $=\left\lvert f(t[i])-f(t[j])\right\rvert\leq L_{f}\left\lvert t[i]-t[j]\right\rvert=L_{f}\Delta t\left\lvert i-j\right\rvert$.
The following proposition quantifies how linear constraints transform under coarsening.
@@
Let $a\in\mathbb{R}^{I}$ discretize a bounded $L_{g}$-Lipschitz function $g:[\ell,u]\to\mathbb{R}$, and let $x$ discretize a bounded $L_{f}$-Lipschitz function $f$. Define the Lipschitz constant for the product function $fg$ as $L_{fg}=L_{f}\left\lVert g\right\rVert_{\infty}+L_{g}\left\lVert f\right\rVert_{\infty}$. If $\langle a,x\rangle=b$, then
$\left\lvert\left\langle\overline{a},\overline{x}\right\rangle-\tfrac{1}{2}b\right\rvert$ $\leq\tfrac{1}{4}L_{fg}(u-l)\tfrac{I}{I-1}$ $\quad\text{for even $I$},$
$\left\lvert\left\langle\overline{a},\overline{x}\right\rangle-\tfrac{1}{2}\tfrac{I+1}{I}b\right\rvert$ $\leq\tfrac{1}{2}L_{fg}(u-l)$ $\quad\text{for odd $I$}.$
Proof 3.5.
Proposition˜3.4 extends naturally to multiple linear constraints, as described in the following corollary.
@@
Let $g_{k}:[\ell,u]\to\mathbb{R}$ be bounded $L_{g_{k}}$-Lipschitz functions for $k\in[K]$. Let $A\in\mathbb{R}^{K\times I}$ be the discretization $A[k,i]=g_{k}(t[i])$, and define $\overline{A}=A[:,\texttt{begin}\mathbin{:}2\mathbin{:}\texttt{end}]$ as the $(K\times\left\lceil I\right\rceil_{e}/2)$-submatrix with every other column removed, where $\left\lceil I\right\rceil_{e}$ rounds up to the nearest even integer. Define $L_{fg}=\max_{k\in[K]}(L_{fg_{k}})=\max_{k\in[K]}(L_{f}\left\lVert g_{k}\right\rVert_{\infty}+L_{g_{k}}\left\lVert f\right\rVert_{\infty})$ as the maximum Lipschitz constant of the product functions $\{fg_{k}\}_{k\in[K]}$. If $Ax=b\in\mathbb{R}^{K}$, then
$\left\lVert\overline{A}\overline{x}-\tfrac{1}{2}b\right\rVert_{2}$ $\leq\sqrt{K}\cdot\tfrac{1}{4}L_{fg}(u-l)\tfrac{I}{I-1}$ $\quad\text{for even $I$},$
$\left\lVert\overline{A}\overline{x}-\tfrac{I+1}{I}\tfrac{1}{2}b\right\rVert_{2}$ $\leq\sqrt{K}\cdot\tfrac{1}{2}L_{fg}(u-l)$ $\quad\text{for odd $I$}.$
Proof 3.7.
Let $a_{k}=A[k,:]\in\mathbb{R}^{I}$ be the $k$th row vector in $A$. By Proposition˜3.4,
$\textstyle\left\lVert\overline{A}\overline{x}-\frac{b}{2}\right\rVert_{2}^{2}$ $=\textstyle\sum_{k=1}^{K}\left\lvert\left\langle\overline{a}_{k},\overline{x}\right\rangle-\frac{b[k]}{2}\right\rvert^{2}\leq\textstyle\sum_{k=1}^{K}\left(\tfrac{I}{I-1}\tfrac{L_{fg_{k}}(u-l)}{4}\right)^{2}$
$\leq\left(\tfrac{I}{I-1}\tfrac{u-l}{4}\right)^{2}\textstyle\sum_{k=1}^{K}\max_{k}\{(L_{fg_{k}})^{2}\}=\left(\tfrac{I}{I-1}\tfrac{u-l}{4}\right)^{2}(L_{fg})^{2}K,$
for even $I$. Taking square roots completes this case, and the case for odd $I$ is similar.
@@
$\left\lvert\left\lVert\overline{x}\right\rVert_{1}-\tfrac{1}{2}b\right\rvert$ $\leq\tfrac{1}{4}L_{f}(u-l)\tfrac{I}{I-1}$ $\quad\text{for even $I$},$
$\left\lvert\left\lVert\overline{x}\right\rVert_{1}-\tfrac{I+1}{I}\tfrac{1}{2}b\right\rvert$ $\leq\tfrac{1}{2}L_{f}(u-l)$ $\quad\text{for odd $I$}.$
Proof 3.9.
The proof is similar to the proof of Proposition˜3.4 (see Section˜A.1) by considering the vector $a=\mathrm{sign}(x)$, and using the Lipschitz constant $L_{f}$ for $f$ to bound $\left\lvert x[i]-x[i+1]\right\rvert$ in place of $L_{fg}$.
@@
$\left\lvert\left\lVert\overline{z}\right\rVert_{1}-\tfrac{1}{2}\right\rvert\leq\tfrac{1}{4b}L_{f}(u-l)\tfrac{I}{I-1}.$
@@ 4 Convergence of a Multiscale Method
Both the standard and multiscale approaches converge because, by design, they both apply a convergent algorithm at the finest scale. The multiscale approach achieves faster convergence by warm starting the fine-scale iterations with solutions from coarser scales. The next subsections develop this analysis.
@@ 4.1 Functional Analysis Lemmas
Figure 5: Example for Lemma˜4.1 with the function $x=f(t)=t-\cos(3\pi t)$ on the interval $\ell,u]=[0,1]$. The function $f$ must lie within the dashed parallelogram since it is $L_{f}=1+3\pi$ Lipschitz. Lemma˜4.1 uses the parallelogram constraint to bound the distance between $f$ and the linear interpolation $\lambda_{1}f(a)+\lambda_{2}f(b)$. Lemma˜4.3 shows that the bound in Eq.˜5 is tight for any given $\lambda_{1},\lambda_{2}\geq 0$ with $\lambda_{1}+\lambda_{2}=1$.
@@
Given $\lambda_{1},\lambda_{2}\geq 0$ with $\lambda_{1}+\lambda_{2}=1$, an interval $\ell,u]$, and constant $L_{f}>0$, there exists an $L_{f}$-Lipschitz function on $\ell,u]$ such that Eq.˜5 holds with equality.
@@
Construct $f(t)=L_{f}\left\lvert t-(\lambda_{1}\ell+\lambda_{2}u)\right\rvert$. This function is $L_{f}$-Lipschitz with end-point values $f(\ell)=L_{f}\lambda_{2}\left\lvert\ell-u\right\rvert$ and $f(u)=L_{f}\lambda_{1}\left\lvert\ell-u\right\rvert$, and $f(\lambda_{1}\ell+\lambda_{2}u)=0$. Therefore $\left\lvert\lambda_{1}f(\ell)+\lambda_{2}f(u)-f(\lambda_{1}\ell+\lambda_{2}u)\right\rvert=2L_{f}\lambda_{1}\lambda_{2}\left\lvert\ell-u\right\rvert$.
To bound the error introduced when interpolating a coarse-scale solution to a finer scale, we apply Lemma˜4.1 repeatedly with centre point interpolation ($\lambda_{1}=\lambda_{2}=1/2$). This bounds the error between an exact fine-scale discretization $x_{s}^{*}[j]=f(t_{s}[j])$ and the linear interpolation $x_{s}=\underline{x}_{s+1}$ obtained from a coarser discretization $x_{s+1}[i]=f(t_{s+1}[i])$. See Fig.˜6 (top).
@@
Let $f:[\ell,u]\to\mathbb{R}$ be $L_{f}$-Lipschitz continuous. On the interval $\ell,u]$, let $t_{s}$ be a fine grid with J=2I1J=2I-1 points, and $t_{s+1}$ be a coarse grid with $I$ points. Discretize the function exactly at both scales to obtain vectors $x^{*}_{s}\in\mathbb{R}^{J}$ and $x^{*}_{s+1}\in\mathbb{R}^{I}$ where $x^{*}_{s}[j]=f(t_{s}[j])$ and $x^{*}_{s+1}[i]=f(t_{s+1}[i])$. Let $x_{s+1}=x^{*}_{s+1}\in\mathbb{R}^{I}$ and linearly interpolate to obtain $x_{s}=\underline{x}_{s+1}\in\mathbb{R}^{J}$ (see Definition˜2.2). Then the difference between the interpolated $x_{s}$ and exact values $x_{s}^{*}$ is bounded by
(6) $\left\lVert x_{s}-x_{s}^{*}\right\rVert_{2}\leq L_{f}\tfrac{1}{2\sqrt{I-1}}\left\lvert u-\ell\right\rvert$.}
@@
When the coarse discretization contains error, the interpolation analysis extends naturally. Suppose we interpolate not from exact values $x_{s+1}[i]=x_{s+1}^{*}[i]=f(t_{s+1}[i])$, but from approximate values $x_{s+1}[i]=x_{s+1}^{*}[i]+\delta[i]=f(t_{s+1}[i])+\delta[i]$, where $\delta$ is an $I$-vector that represents the error at the coarse scale. Fig.˜6 (bottom) visualizes this scenario.
@@
(7) $\left\lVert x_{s}-x_{s}^{*}\right\rVert_{2}\leq\sqrt{2}\left\lVert\delta\right\rVert_{2}+L_{f}\tfrac{1}{2\sqrt{I-1}}\left\lvert u-\ell\right\rvert$.}
@@
Figure 6: Illustration of Lemma˜4.5 (left) and Lemma˜4.7 (right) for the function $f(t)=72t^{3}-80t^{2}+10t+1$. Both lemmas bound the error between the true fine-scale discretization $x^{*}_{s}$ (orange circles) and the linear interpolation $x_{s}=\underline{x}_{s+1}$ (green squares) obtained from a coarser-scale discretization. The left figure shows Eq.˜6, where $x_{s+1}=x^{*}_{s+1}$ is exact. The right figure shows Eq.˜7, where $x_{s+1}=x^{*}_{s+1}+\delta$ contains error.
@@ 4.2 Greedy Multiscale
We use the following approach to show convergence of the greedy multiscale algorithm described in Algorithm˜1. At each scale $s$, we use the iterate convergence (Definition˜2.6) to bound the error between our iterate $x_{s}^{K_{s}}$ and the solution $x_{s}^{*}$ at that scale. Then we use the inexact interpolation lemma (Lemma˜4.7) to bound the error between our iterate $x_{s}^{K_{s}}$ and interpolation $\underline{x}_{S}^{K_{S}}$. This interpolated iterate becomes the initialization $x_{S-1}^{0}=\underline{x}_{S}^{K_{S}}$ at the next finest scale.
@@
Let $q$ denote the rate of convergence of an algorithm applied to a loss function $\tilde{\mathcal{L}}$, let $L_{f^{*}}$ be the Lipschitz constant of the solution function $f^{*}$, and let $S$ denote the coarsest scale. The greedy multiscale method in Algorithm˜1 returns iterate $x_{1}^{K_{1}}$ at the finest scale satisfying
$\left\lVert x^{K_{1}}_{1}-x^{*}_{1}\right\rVert_{2}\leq\sqrt{2^{S-1}}(q)^{r_{S}}\left\lVert x_{S}^{0}-x_{S}^{*}\right\rVert_{2}+L_{f^{*}}\tfrac{\left\lvert u-\ell\right\rvert}{2\sqrt{2^{S+1}}}\sum\nolimits_{s=1}^{S-1}2^{s}(q)^{r_{s}},$